OpenStudy (anonymous):

Lilly measures the volume of a sandbox to be 12,960 cubic inches. The actual volume is 11,484 cubic inches. What is the perent error to the nearest tenth of percent, in calculating the volume?

OpenStudy (radar):

I am not sure I remember how this is done but here is something to think about. The error was 12960-11484 = 1476 cu in. (positive) this error would be 1476/11484 or 0.1285 converting to percent would be 12.85%

OpenStudy (anonymous):

okay, sounds reasonable

OpenStudy (anonymous):

The percent error formula is \[\left| your calculation - actual value \right| / actual value \times 100 %\]

OpenStudy (anonymous):

Radar is correct, but there's the formula.

OpenStudy (radar):

Thanks, I was wondering if it was the actual value or the calculated value that the error was divided by, glad to see my instincts were right. lol

OpenStudy (anonymous):

No, the top is the absolute value of the two.

OpenStudy (radar):

Ok, then, I was thinking the sign could indicate the direction of the error, but if it is absolute then it would be irrelevant.

OpenStudy (anonymous):

THANKS FOR ALL YORU HELP!